# VLM-Assisted Continual learning for Visual Question Answering in Self-Driving

### v1

@@ Abstract
In this paper, we propose a novel approach for solving the Visual Question Answering (VQA) task in autonomous driving by integrating Vision-Language Models (VLMs) with continual learning. In autonomous driving, VQA plays a vital role in enabling the system to understand and reason about its surroundings. However, traditional models often struggle with catastrophic forgetting when sequentially exposed to new driving tasks, such as perception, prediction, and planning, each requiring different forms of knowledge. To address this challenge, we present a novel continual learning framework that combines VLMs with selective memory replay and knowledge distillation, reinforced by task-specific projection layer regularization. The knowledge distillation allows a previously trained model to act as a ”teacher” to guide the model through subsequent tasks, minimizing forgetting. Meanwhile, task-specific projection layers calculate the loss based on the divergence of feature representations, ensuring continuity in learning and reducing the shift between tasks. Evaluated on the DriveLM dataset, our framework shows substantial performance improvements, with gains ranging from $21.40\%$ to $32.28\%$ across various metrics. These results highlight the effectiveness of combining continual learning with VLMs in enhancing the resilience and reliability of VQA systems in autonomous driving. We will release our source code.
@@ 1 Introduction
With the rapid advancement of technology, autonomous driving [30, 38] has become a landmark achievement in the development of modern transportation systems. Self-driving cars utilize various advanced sensors, cameras, and artificial intelligence algorithms to interpret the traffic environment, promising a safer and more efficient driving experience. Nowadays, Vision-language models [11, 31, 25, 16, 19], play a crucial role as the core of autonomous driving systems. These models understand and predict road conditions and their complexity by parsing visual images and video data collected from cameras and other sensors, combined with simultaneously processed verbal commands.
@@
Although vision-language models (VLMs) [11, 31] show great potential in handling complex driving tasks, these models still face multiple challenges in autonomous driving applications. With the diversity of task types including perception, prediction, and planning, VLMs in autonomous driving increasingly need to be equipped with strong continual learning capabilities. Continual learning [4, 14, 28, 27, 37] allows models to accumulate new knowledge based on existing knowledge without having to go back to initial training data to relearn. Currently, continual learning is also introduced in a number of VLM models and VQA models in mitigating catastrophic forgetting [37, 15] in different domains, e.g., medical domain [35, 2, 7, 33] and so on. However, in the context of autonomous driving, the application of these methods requires special adaptation and optimization. Any learning inaccuracies may lead to serious safety consequences. Addressing catastrophic forgetting within the VLM framework for autonomous driving tasks—such as perception, prediction, and planning—requires innovative strategies beyond standard sequential training. As illustrated in Figure 1, VLMs trained on new tasks may experience catastrophic forgetting, where prior task knowledge is overwritten, leading to loss of critical information. Therefore, ensuring that VLMs are able to adapt to new driving environments and tasks without losing prior knowledge is crucial for improving the reliability and safety of autonomous driving systems.
@@
To address above-mentioned issues, in this work, we introduce a novel hybrid continual learning strategy with VLMs, which combines selective memory replay [5] with knowledge distillation [23], enhanced by projection layer regularization. Specially, selective memory replay periodically reinforces previously learned information to mitigate forgetting, while knowledge distillation enables the model trained on earlier tasks to act as a teacher, guiding subsequent tasks and preserving complex task knowledge. This strategy balances learning new tasks and retaining essential prior knowledge, crucial in a continually changing autonomous driving environment. By replaying real driving data, the system retains critical safety-related information across diverse driving conditions, thus enhancing the model’s reliability and generalization capability.
@@
To further strengthen our model against the challenge of catastrophic forgetting, we introduce a new task-specific regularization methods of projection layers. In models with a relatively large number of model parameters and experimental data, both the classical continual learning parameter regularization approach [14] and the gradient projection approach [20] face challenges in terms of computational resource requirements and operational efficiency. Therefore, following the idea of regularization, we regularize the learning process by carefully designing the model feature projection layer and calculating the loss based on the differences in feature representations in successive tasks. This regularization not only reduces additional computational and memory requirements, but also prevents the dilution of previously learned features and ensures that the model’s learning trajectory closely follows the continuum, thus mitigating the effects of sudden knowledge transfer.
@@
    We benchmark our model on the DriveLM dataset, achieving significant performance improvements. Compared to the baseline model, our method demonstrated improvements across various metrics, ranging from a minimum of $21.40\%$ to a maximum of $32.28\%$, highlighting its superior capacity to handle complex multimodal VQA tasks in autonomous driving environments.
@@ 2 Related Work
Vision Language Models for VQA. In the field of autonomous driving, the application of Vision-Language Models (VLMs) is rapidly expanding [34, 31], enhancing the system’s understanding of and decision-making capabilities in complex driving environments. VLMs, through extensive image-text pre-training, have already provided the ability for zero-shot learning in autonomous driving [36]. For example, the DriveVLM [31] system explores the integration of VLMs within traditional autonomous driving technologies, enhancing the vehicle’s spatial reasoning and planning capabilities. Additionally, the DriVLMe [13] and Co-driver [12] projects explore how LLMs mimic human understanding and behavior in handling complex road conditions from the perspectives of experiential learning and social interaction, respectively. In recent years, Vision-Question-Answering (VQA) datasets for autonomous driving, such as the Nuscenes-QA [24]dataset, DriveLM dataset [29], have further facilitated advancements in the multimodal understanding of driving scenarios. In our model, we delve deeper into the field of continual learning for autonomous driving, building upon the EM-VLM4AD [11] framework.
Continual Learning. Continual learning (CL) techniques like memory replay [4], regularization [14, 1], and optimization [20, 9] mitigate catastrophic forgetting by balancing new and old tasks. Memory replay stores key data or features in a buffer, using methods such as random selection [6] or feature averaging [26]. Regularization approaches add terms that selectively preserve crucial parameters based on their importance, assessed using tools like the Fisher Information Matrix (FIM) [14, 28]. Optimization methods, such as OGD [9], maintain previous gradient directions while adjusting current gradients for orthogonality. However, in our approach, we employ some techniques for text vectorization and clustering of different tasks in the dataset to optimize data selection in the memory replay mechanism. In addition to this, we naturally combine memory replay with knowledge distillation (KD), which additionally integrates the past information of the old model. Due to the relatively large number of parameters in the VLM, regularization with gradient projection via parametric aspects is not very practical. Therefore, we emulated the idea of regularization and projection. We did this by comparing the projection layer gaps between different tasks after the model’s feature embedding was introduced. Additionally, we added extra terms to the loss function to further enhance the continual learning effect at the model’s feature level.
@@ 3.1 Overview
Targeting the autonomous driving visual question answering (VQA) dataset, we propose a new continual learning approach for Vision-Language Model (VLMs). This method alternates continual learning in two phases: one part is carried out during the data replay phase, and the other part during the current task data training phase. We construct a simple autonomous driving VLM using a straightforward image embedding network (ViT) [8] combined with a pre-trained T5 language model, applying the designed continual learning method to this VLM. The visualization of our continual learning model is shown in Figure 1, which includes memory reply with knowledge distillation and embedding projection. Next, we will delve into the details of each continual learning phase.
3.2 Memory Reply with Knowledge Distillation
@@
Memory Reply. Firstly, we perform text vectorization on the data in the task and use TF-IDF to extract features from the problem set, which helps us understand the main semantic content of the data. Then, by applying the K-means algorithm, we cluster the questions into several categories, which ensures that the data stored in the memory is representative and diverse in terms of contents. For each clustering, we refine the selection process: the number of samples to be selected for each category, denoted by $S_{c,n}$, is determined based on the proportion of data in each category, ensuring that the importance of each category is balanced. Specifically, we calculate the ratio of the total data volume $D_{n}$ to the data volume $D_{c,n}$ in each category $c$ within task $D_{n}$ , and use this ratio to determine $S_{c,n}$. This approach not only improves the coverage of the samples but also increases the adaptability and robustness of the model to different types of problems. $S_{c,n}$ can be formulated as the following:
@@
Knowledge Distillation. In our approach, knowledge distillation guides the current model $M_{s}$ by utilizing the model trained in the previous task as the teacher model $M_{t}$. The teacher model provides insight into the specific details of the previous task and outputs the corresponding output probabilities encoded as $K_{t}$. These probabilities were adjusted by temperature parameters to enhance the smoothness of the distribution, making subtle decision boundaries more easily captured by the student model. The ability of the student model to replicate the teacher’s decision-making process was effectively quantified by calculating the loss of the teacher model from the current model’s predictions probabilities $K_{s}$ through KL scatter during training, which can be formulated as the following:
$\mathcal{L}_{KD}(K_{t}\parallel K_{s})=\sum_{i}K_{t}(i)\log\left(\frac{K_{t}(i)}{K_{s}(i)}\right),$ (3)
where $K_{t}(i)$ and $K_{s}(i)$ represent output probabilities of the teacher and student models for category $i$, respectively.
By adjusting the balance between the distillation loss $\mathcal{L}_{KD}$ and the traditional task-specific loss $\mathcal{L}_{GT}$ in the autonomous driving task, we filtered the replay data according to the data characteristics of the dataset to ensure as much comprehensiveness of the data from past tasks as possible. At the same time, knowledge distillation was introduced to optimize the benefits between direct task learning and the detailed guidance provided by the teacher model. Performance can be improved by mixing losses to ensure that the model is not only capable of basic continual learning properties, but can also be adapted to new tasks in terms of understanding complex autonomous driving interactions. The loss function for the network in memory replay is formulated as
$\mathcal{L}oss=\alpha\mathcal{L}_{KD}+(1-\alpha)\mathcal{L}_{GT},$ (4)
where $\alpha$ is the parameter controlling the weight between knowledge distillation loss and task-specific loss.
@@ 3.3 Embedding Projection
In the continual learning task of autonomous driving, while memory replay and knowledge distillation can guide and assist the model in learning more abstract features and decision logic, this approach relies on the accuracy and generalization ability of the teacher’s model with the replayed data. Furthermore, it primarily targets the similarity of the model outputs rather than directly addressing the feature continuity across tasks. Accordingly, the considerable bias of the model resulting from training on new tasks can be constrained and mitigated by regularizing the model during the phase when the model is trained on the current data for each task. In the context of the autonomous driving VLM model, the parameter regularization and gradient projection techniques that are typically applicable to small-scale networks cannot be directly applied. Consequently, our approach involves introducing a projection layer into the feature embedding block of the model(Figure 2(b)). By updating the feature projection layer and regularizing the model during the training of each task, it is possible to maintain the overall structure of the model’s features without distorting them by new tasks.
@@
where $\mathbf{e}$ represents the feature embeddings obtained from the multimodal input processing. The transformation is designed to ensure that each task’s feature representation is both unique and aligned with the task-specific requirements.
@@
$\mathcal{L}oss_{total}=\begin{cases}\alpha\mathcal{L}_{\text{KD}}+(1-\alpha)\mathcal{L}_{\text{GT}}&\text{if replay data},\\ \lambda\mathcal{L}_{{pro}}+\mathcal{L}_{\text{GT}}&\text{otherwise}.\end{cases}$ (9)
@@ 4.1 Experimental Settings
Dataset. We use the DriveLM [29] dataset, which is a multimodal dataset designed for autonomous driving research, containing a large amount of image and text data to support the development of vision-language models. In our experiments, we divided the DriveLM dataset into 4 tasks: Perception, Prediction, Planning, and Behavior, and each task contains a corresponding training/validation/testing set. More details please refer to supplementary documents.
Metrics. In this experiment, we use a variety of evaluation metrics to comprehensively assess the performance of the model. Specifically, BLEU [21], which is used to evaluate the fluency and accuracy of the generated text; METEOR [3], which takes into account synonyms and sentence structure which is used to assess translation quality; ROUGE [18], which is mainly used to assess the degree of overlap between automatic digest or machine translation outputs and human referencing; and CIDEr [32], which is specifically used to assess the quality of image description. These metrics assess the model’s capabilities in language generation and comprehension for critical autonomous driving tasks: Perception, Prediction, Planning, and Behavior. The dataset includes task-specific questions in a Visual Question Answering (VQA) format to evaluate the performance of the model, ensuring alignment with real-world autonomous driving needs.
Method BLUE-1 BLUE-2 BLUE-3 BLUE-4 METEOR ROUGE_L CIDEr
@@
ER KD PRO BLUE-4 METEOR ROUGE_L CIDEr

### v2

@@ Abstract
In this paper, we propose a novel approach for solving the Visual Question Answering (VQA) task in autonomous driving by integrating Vision-Language Models (VLMs) with continual learning. In autonomous driving, VQA plays a vital role in enabling the system to understand and reason about its surroundings. However, traditional models often struggle with catastrophic forgetting when sequentially exposed to new driving tasks, such as perception, prediction, and planning, each requiring different forms of knowledge. To address this challenge, we present a novel continual learning framework that combines VLMs with selective memory replay and knowledge distillation, reinforced by task-specific projection layer regularization. The knowledge distillation allows a previously trained model to act as a ”teacher” to guide the model through subsequent tasks, minimizing forgetting. Meanwhile, task-specific projection layers calculate the loss based on the divergence of feature representations, ensuring continuity in learning and reducing the shift between tasks. Evaluated on the DriveLM dataset, our framework shows substantial performance improvements, with gains ranging from $20.11\%$ to $35.16\%$ across various metrics. These results highlight the effectiveness of combining continual learning with VLMs in enhancing the resilience and reliability of VQA systems in autonomous driving. We will release our source code.
@@ 1 Introduction
With the rapid advancement of technology, autonomous driving [aitech, autonomoustechnology] has become a landmark achievement in the development of modern transportation systems. Self-driving cars utilize various advanced sensors, cameras, and artificial intelligence algorithms to interpret the traffic environment, promising a safer and more efficient driving experience. Nowadays, Vision-language models [EM-VLM4AD, Drivevlm, clip, blip, llava], play a crucial role as the core of autonomous driving systems. These models understand and predict road conditions and their complexity by parsing visual images and video data collected from camera and other sensors, combined with simultaneously processed verbal commands.
@@
Although vision-language models (VLMs) [EM-VLM4AD, Drivevlm] show great potential in handling complex driving tasks, these models still face multiple challenges in autonomous driving applications. With the diversity of task types including perception, prediction, and planning, VLMs in autonomous driving increasingly need to be equipped with strong continual learning capabilities. Continual learning [DER, EWC, EWC++, MER, vqacl] allows models to accumulate new knowledge based on existing knowledge without having to go back to initial training data to relearn. Currently, continual learning is also introduced in a number of VLM models and VQA models in mitigating catastrophic forgetting [vqacl, lei2023symbolic] in different domains, e.g., medical domain [yu2024boosting, bai2023revisiting, chen2024llm, yi2023towards] and so on. However, in the context of autonomous driving, the application of these methods requires special adaptation and optimization. Any learning inaccuracies may lead to serious safety consequences. Addressing catastrophic forgetting within the VLM framework for autonomous driving tasks—such as perception, prediction, and planning—requires innovative strategies beyond standard sequential training. As illustrated in Figure 1, VLMs trained on new tasks may experience catastrophic forgetting, where prior task knowledge is overwritten, leading to loss of critical information. Therefore, ensuring that VLMs are able to adapt to new driving environments and tasks without losing prior knowledge is crucial for improving the reliability and safety of autonomous driving systems.
@@
To address above-mentioned issues, in this work, we introduce a novel hybrid continual learning strategy with VLMs, which combines selective memory replay [ER] with knowledge distillation [KD], enhanced by projection layer regularization. Specially, selective memory replay periodically reinforces previously learned information to mitigate forgetting, while knowledge distillation enables the model trained on earlier tasks to act as a teacher, guiding subsequent tasks and preserving complex task knowledge. This strategy balances learning new tasks and retaining essential prior knowledge, crucial in a continually changing autonomous driving environment. By replaying real driving data, the system retains critical safety-related information across diverse driving conditions, thus enhancing the model’s reliability and generalization capability.
@@
To further strengthen our model against the challenge of catastrophic forgetting, we introduce a new task-specific regularization methods of projection layers. In models with a relatively large number of model parameters and experimental data, both the classical continual learning parameter regularization approach [EWC] and the gradient projection approach [GEM] face challenges in terms of computational resource requirements and operational efficiency. Therefore, following the idea of regularization, we regularize the learning process by carefully designing the model feature projection layer and calculating the loss based on the differences in feature representations in successive tasks. This regularization not only reduces additional computational and memory requirements, but also prevents the dilution of previously learned features and ensures that the model’s learning trajectory closely follows the continuum, thus mitigating the effects of sudden knowledge transfer.
@@
    We benchmark our model on the DriveLM dataset, achieving significant performance improvements. Compared to the baseline model, our method demonstrated improvements across various metrics, ranging from a minimum of $20.11\%$ to a maximum of $35.16\%$, highlighting its superior capacity to handle complex multimodal VQA tasks in autonomous driving environments.
@@ 2 Related Work
Vision Language Models for VQA. In the field of autonomous driving, the application of Vision-Language Models (VLMs) is rapidly expanding [you2024v2x, Drivevlm], enhancing the system’s understanding of and decision-making capabilities in complex driving environments. VLMs, through extensive image-text pre-training, have already provided the ability for zero-shot learning in autonomous driving [vlmsurvey]. For example, the DriveVLM [Drivevlm] system explores the integration of VLMs within traditional autonomous driving technologies, enhancing the vehicle’s spatial reasoning and planning capabilities. Additionally, the DriVLMe [drivlme] and Co-driver [Co-driver] projects explore how LLMs mimic human understanding and behavior in handling complex road conditions from the perspectives of experiential learning and social interaction, respectively. In recent years, Vision-Question-Answering (VQA) datasets for autonomous driving, such as the Nuscenes-QA [nuscenesqa]dataset, DriveLM dataset [drivelm], have further facilitated advancements in the multimodal understanding of driving scenarios. Meanwhile, topology-aware reasoning has also emerged as an important direction. For instance, T2SG [lv2025t2sg] introduces a traffic-topology scene graph to explicitly model relational structures in driving environments, providing complementary insights for question answering systems. Other efforts enhance driving VLMs through auxiliary predictive tasks, such as uncertainty-guided attention prediction [zhu2023unsupervised] or knowledge-graph-based retrieval augmentation [ye2025safedriverag], both of which improve robustness and interpretability in complex driving scenes. In our model, we delve deeper into the field of continual learning for autonomous driving, building upon the EM-VLM4AD [EM-VLM4AD] framework.
Continual Learning. Continual learning (CL) techniques like memory replay [DER], regularization [EWC, MAS], and optimization [GEM, OGD] mitigate catastrophic forgetting by balancing new and old tasks. Memory replay stores key data or features in a buffer, using methods such as random selection [chaudhry2019tiny] or feature averaging [iCaRL]. Regularization approaches add terms that selectively preserve crucial parameters based on their importance, assessed using tools like the Fisher Information Matrix (FIM) [EWC, EWC++]. Optimization methods, such as OGD [OGD], maintain previous gradient directions while adjusting current gradients for orthogonality. However, in our approach, we employ some techniques for text vectorization and clustering of different tasks in the dataset to optimize data selection in the memory replay mechanism. In addition to this, we naturally combine memory replay with knowledge distillation (KD), which additionally integrates the past information of the old model. Due to the relatively large number of parameters in the VLM, regularization with gradient projection via parametric aspects is not very practical. Therefore, we emulated the idea of regularization and projection. We did this by comparing the projection layer gaps between different tasks after the model’s feature embedding was introduced. Additionally, we added extra terms to the loss function to further enhance the continual learning effect at the model’s feature level.
@@ 3.1 Overview
Targeting the autonomous driving visual question answering (VQA) dataset, we propose a new continual learning approach for Vision-Language Model (VLMs). This method alternates continual learning in two phases: one part is carried out during the data replay phase, and the other part during the current task data training phase. We construct a simple autonomous driving VLM using a straightforward image embedding network (ViT) [ViT] combined with a pre-trained T5 language model, applying the designed continual learning method to this VLM. The visualization of our continual learning model is shown in Figure 1, which includes memory replay with knowledge distillation and embedding projection. This unified design offers a coordinated solution to the stability–plasticity trade-off in continual multimodal learning, achieving complementary effects unattainable with isolated strategies. Next, we will delve into the details of each continual learning phase.
3.2 Memory replay with Knowledge Distillation
@@
Memory replay. Firstly, we perform text vectorization on the data in the task and use TF-IDF to extract features from the problem set, which helps us understand the main semantic content of the data. Then, by applying the K-means algorithm, we cluster the questions into several categories, which ensures that the data stored in the memory is representative and diverse in terms of contents. For each clustering, we refine the selection process: the number of samples to be selected for each category, denoted by $S_{c,n}$, is determined based on the proportion of data in each category, ensuring that the importance of each category is balanced. Specifically, we calculate the ratio of the total data volume $D_{n}$ to the data volume $D_{c,n}$ in each category $c$ within task $D_{n}$ , and use this ratio to determine $S_{c,n}$. This approach not only improves the coverage of the samples but also increases the adaptability and robustness of the model to different types of problems. $S_{c,n}$ can be formulated as the following:
@@
Knowledge Distillation. In our approach, the previous task model $M_{t}$ serves as the “teacher” and the current model $M_{s}$ as the “student.” Rather than distilling all tokens equally, we only apply soft‐label distillation to those tokens for which the teacher is sufficiently confident; for less certain tokens we rely solely on ground‐truth supervision. Given the teacher’s logits $z^{T}_{i,t}$ for sample $i$ at token position $t$, we first compute temperature‐scaled probabilities. Here, $z^{T}_{i,t}\in\mathbb{R}^{K}$ denotes the $K$-dimensional logit vector over the vocabulary.
$p^{T}_{i,t}=\operatorname{softmax}\bigl(z^{T}_{i,t}/T\bigr)\,$, (3)
where $T$ is the distillation temperature. We then measure the token‐level confidence by
$c_{i,t}=\max_{k}p^{T}_{i,t}[k]\,$, (4)
and map it to a per‐token distillation weight
$\alpha_{i,t}=\begin{cases}0,&c_{i,t}<\tau,\\[6.0pt] \displaystyle\frac{c_{i,t}-\tau}{1-\tau}\,\alpha_{\max},&c_{i,t}\geq\tau,\end{cases}$ (5)
where $\tau$ is the confidence threshold and $\alpha_{\max}$ the maximum distillation ratio.
Similarly, the student’s temperature‐scaled probabilities are
$p^{S}_{i,t}=\operatorname{softmax}\bigl(z^{S}_{i,t}/T\bigr)\,.$ (6)
For each valid (non‐padding) token we compute the standard cross‐entropy loss $\mathcal{L}_{GT}^{(i,t)}$ using the ground‐truth label, and the KL‐divergence
$\mathcal{L}_{KD}^{(i,t)}=\sum_{k}p^{T}_{i,t}[k]\,\log\frac{p^{T}_{i,t}[k]}{p^{S}_{i,t}[k]}\times T^{2}$ (7)
for soft‐label alignment. The overall loss for a batch is then
$\mathcal{L}_{\text{replay}}=\frac{1}{\sum_{i,t}m_{i,t}}\sum_{i,t}m_{i,t}\Bigl[(1-\alpha_{i,t})\,\mathcal{L}_{GT}^{(i,t)}+\alpha_{i,t}\,\mathcal{L}_{KD}^{(i,t)}\Bigr],$ (8)
where $m_{i,t}\in\{0,1\}$ masks out padding tokens. This selective distillation ensures that the student inherits only the teacher’s highly reliable “dark knowledge,” while avoiding the teacher’s uncertain or incorrect predictions, thus striking a better balance between retaining past‐task performance and adapting to new tasks.
@@ 3.3 Embedding Projection
In the continual learning setting of autonomous driving, memory replay and knowledge distillation help the model learn abstract features and decision logic. However, their effectiveness depends on the teacher model’s accuracy and generalization over replayed data. Moreover, these methods focus on output similarity, rather than directly preserving feature continuity across tasks. To mitigate the bias caused by training on new tasks, we introduce a model regularization phase during each task’s learning process. In the context of large-scale VLMs for autonomous driving, traditional parameter regularization and gradient projection methods used in smaller networks are not directly applicable. Thus, we propose incorporating a feature projection layer into the model’s embedding block (Figure 2(b)). By updating this projection and applying feature-level regularization during training, we maintain the structural consistency of the model’s representations across tasks.
@@
where $\mathbf{e}$ represents the feature embeddings obtained from the multimodal input processing. The transformation is designed to ensure that each task’s feature representation is both unique and aligned with the task-specific requirements. It is important to note that the projection layers do not freeze or isolate the underlying embedding $\mathbf{e}$. Instead, each task learns its own projection $P_{n}$ that maps $\mathbf{e}$ into a lower-dimensional, task-specific space. This design serves two purposes. First, by constraining the similarity of the projected features rather than the raw embeddings, we reduce the variance introduced by task-dependent input distributions, making the feature space more stable across tasks. Second, even if $\mathbf{e}$ evolves during training, the projection spaces act as an anchoring mechanism: the task-specific representations Eq. (9) remain aligned through the regularization, effectively controlling representational drift without over-constraining the backbone VLM.
Algorithm 1 continual learning for AD VLM
Input
Output
Trained VLM model $\phi$
1: Initialize $\phi$ and $R$
2:for $n=1$ to NN do
3:  Load task-specific dataset $D_{n}$ and task $T_{n}$
4:  if $n==1$ then
5:   Train $\phi$ on $D_{1}$
6:   Populate $R$ using TF-IDF and K-means on $D_{1}$
7:  else
8:   for each iteration do
9:    Sample $x$ from $D_{n}\cup R$
10:    if $x\in D_{n}$ then
11:     Update $\phi$ using projection layer $P_{n}$
12:    else
13:     Replay using model $M_{t}$ from $T_{n-1}$
14:    end if
15:   end for
16:   Refresh $R$ with samples from $D_{n}$
17:  end if
18:  Evaluate $\phi$ on $T_{n}$
19:end for
@@
$\mathcal{L}oss_{total}=\begin{cases}\alpha\mathcal{L}_{\text{replay}}&\text{if replay data},\\ \lambda\mathcal{L}_{{pro}}+\mathcal{L}_{\text{GT}}&\text{otherwise}.\end{cases}$ (13)
The entire continual learning workflow is summarized in Algorithm 1.
@@ 4.1 Experimental Settings
Method BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE_L CIDEr
@@
Dataset. We use the DriveLM [drivelm] dataset, a multimodal dataset designed for autonomous driving research, containing large amounts of image and text data across various conditions—including sunny, rainy, and nighttime scenarios—to support the development of vision‑language models.In our experiments, we divided the DriveLM dataset into 4 tasks: Perception, Prediction, Planning, and Behavior, and each task contains a corresponding training/validation/testing set. More details please refer to supplementary documents.
Metrics. In our experiment, we adopt a variety of evaluation metrics to comprehensively assess model performance. Specifically, BLEU[Bleu], METEOR[Meteor], ROUGE[Rouge], and CIDEr[Cider] are used following the official DriveLM evaluation protocol. Although these metrics originate from natural-language generation, DriveLM assess task correctness rather than linguistic fluency. This is because the dataset expresses multiple autonomous-driving subtasks in QA form: perception questions correspond to object detection and attribute recognition; prediction questions describe the future motion of surrounding agents; planning questions encode rule-based driving decisions; and behavior questions reflect high-level driving intent. Thus, the textual answers function as structured labels for these subtasks, and the metric scores reflect how accurately the model solves Perception, Prediction, Planning, and Behavior tasks.
Under our continual learning setup, after training each task $T_{n}$, the model is re-evaluated on all previously learned tasks. As a result, questions from early tasks may accumulate multiple predictions across training stages. In line with standard continual-learning practice, after completing all tasks, we report the scores of the final model on the test sets of every task.
4.2 Analysis
Compared Methods. We compare our method with the following methods on the DriveLM dataset [drivelm]: the weight regularization method EWC [EWC], the function regularization method LwF [LwF], and the memory replay method DER [DER] to validate the effectiveness of our approach. In addition, we provide a lower bound (Vanilla), which simply performs gradient updating without any countermeasures against forgetting, and a relative upper bound (Joint), which trains all tasks jointly. Moreover, we compare another method that uses a joint training approach but is based on a different model, referred to as DriveLM-Agent [drivelm], to demonstrate the performance differences in handling multitask learning across various model architectures.
Implementation Details. Our implementation is based on the EM‑VLM4AD framework [EM-VLM4AD], utilizing a Vision Transformer (ViT)[ViT] for multi-view image embedding and a pre-trained T5 model for text processing, enabling Visual Question Answering (VQA) in autonomous driving. Our EM‑VLM4AD (T5‑Base) model is lightweight, with 235M parameters, 9.47B FLOPs, and 0.94GB memory, achieving under 1s latency per sample—suitable for near real‑time use. We implement our model in PyTorch[pytorch] and train all tasks on a single NVIDIA L20 GPU, with each task trained for 4 epochs. In our four-task setting, the complete continual learning pipeline takes about 20 hours to finish, and the additional overhead introduced by replay (small batch fraction), knowledge distillation (one extra teacher forward per replay batch), and projection layers (<<1M parameters, <<0.5% FLOPs) remains lightweight relative to the VLM backbone. For hyper‑parameters, we use a learning rate of 1e‑4, a weight decay of 0.05, a batch size of 4, a temperature parameter set to 2 and an $\alpha$ value of 0.7 for the knowledge distillation part. A dynamic weight parameter adjusts the projection regularization loss, where the $\lambda$ parameter is initialized to 0.05 and halved progressively with each task. More details of experimental hyperparameter settings and experiments refer to supplementary materials.
ER KD PRO BLEU-4 METEOR ROUGE_L CIDEr